# Towards Generalizable AI-Generated Image Detection via Image-Adaptive Prompt Learning

### v2

@@ 3.3 Image-Adaptive Prompt Learning
Specifically, the Test-Time Token Tuning starts with generating multiple views $V\in R^{N_{v}\times 224\times 224\times 3}$ of the original image, where $N_{v}$ is the number of views. In particular, the generated views could be divided into a global view and multiple local views. For a global view, we resize the entire image to the input size of CLIP, i.e. $224\times 224$. For local views, we randomly crop $N_{v}$-1 $224\times 224$ sub-images followed by horizontal flipping augmentation. If the image is too small to crop $N_{v}$-1 different sub-images, we randomly crop $N_{v}$-1 smaller square patches and then resize these patches to $224\times 224$. To reduce the noise from randomly generated views, the Confidence Selection is conducted. We select $m$ views with high predicted confidence for subsequent tuning. The predicted confidence $S_{c}^{i}$ is calculated via Eq. 2.
@@
where $z_{i}$ is the logit calculated by the model of the i-th view. $\sigma(.)$ is the Sigmoid function that transform $z_{i}$ into $(0,1)$.
Given the selected views, we tune the test-time adaptive tokens by minimizing entropy, while other parameters are frozen. The objective function mainly includes averaged entropy loss [50] and pointwise entropy loss [48]. For the averaged entropy loss, its purpose is to improve consistency among predictions across the generated views, and it is shown in Eq. 3 and Eq. 4.
@@
Here, $z_{k}$ is the logit calculated by the model of the $k$-th selected view. Some methods [48] adopt pointwise entropy loss to push the decision boundary away from low-entropy regions, its objective function is shown in Eq. 5.
@@
where $p_{k}$ denotes $\sigma(z_{k})$ of k-th selected view. Based on the results in Tab. 5, we ultimately adopt the averaged entropy loss due to its superior performance across all experimental datasets. Additionally, the above tuning process is performed for a total of $T$ steps.
Conditional Information Learner. To capture subtle and varying cues in AI-generated images, we propose the Conditional Information Learner, which extracts conditional information tailored to each input. It selects the most texture-rich region likely to contain forgery artifacts, enabling focused and efficient feature extraction. This emphasizes low-level patterns often overlooked by the high-level semantic features of the pre-trained CLIP backbone. The extracted information is partitioned into forgery-specific and image-specific components, both directly informed by the input image, thereby enabling the model to achieve more dynamic and effective adaptation to its characteristics.
Specifically, for a given input image, we divide it into local patches $P_{l}\in R^{N_{p}\times 32\times 32\times 3}$, where $N_{p}$ is the number of patches. We use the DCT score [61] to measure the texture richness of each patch and select the patch with the highest DCT value for conditional information extraction. After that, the selected patches are input into high-pass filters [16] to obtain the high-frequency pattern. Two CNN-based feature extractors with identical architectures but separate parameters are used to extract the forgery-specific condition $C_{f}$ and the image-specific condition $C_{i}$, respectively. An auxiliary supervision branch is attached after a forgery-specific condition. It consists of a single-layer MLP, enabling the supervised branch to extract features strongly correlated with the presence of forgery, while the unsupervised branch learns more general features related to the image status. Moreover, as the conditional information is extracted from a small patch, this process is both time-efficient and memory-efficient.
After acquiring the conditional information $C_{i}$ and $C_{f}$, we fuse them with the test-time adaptive tokens $A\in\mathbb{R}^{2\times D}$ via the gated mechanism, to obtain the image-adaptive prompt $P\in\mathbb{R}^{2\times D}$ as shown in Eq. 6.
$P=\{\alpha_{f}*C_{f}+A[0,:],\alpha_{i}*C_{i}+A[1,:]\},$ (6)
where $\alpha_{f}$ and $\alpha_{i}$ denote the learnable gating coefficients associated with the forgery-specific and image-specific conditional information, respectively, and $\{\,\cdot\,$$\cdot\,\}$ represents the concatenation operation.
@@
3.4 Optimal View Selection.
Since images generated by different models often vary considerably in resolution and aspect ratio, previous methods typically adopt resizing or center cropping to standardize the input. However, such fixed preprocessing strategies may discard or distort critical artifacts, as the most informative regions for forgery detection differ across images. To address this limitation, we first select confident views during the Test-Time Token Tuning stage and feed them into the tuned image encoder to obtain confidence scores for each view. Finally, we choose the prediction corresponding to the view with the highest confidence as the final output, thereby reducing the risk of losing forgery-related cues and improving robustness against diverse image sources.
@@ 4.1 Datasets and Metrics
We conduct experiments on two widely used datasets, i.e., UniversalFakeDetect Dataset [40] and GenImage Dataset [68], to demonstrate the effectiveness of our framework.
UniversalFakeDetect Dataset. This dataset includes 20 subsets of generated images. Following previous protocol [40, 58], the real and fake images in ProGAN [25] are used for training, while the testing set contains 19 subsets from ProGAN and other 18 different generative models, including CycleGAN [67], BigGAN [4], StyleGAN [26], GauGAN [41], StarGAN [12], DeepFakes [46], SITD [7], SAN [13], CRN [9], IMLE [30], Guided [14], LDM [45], Glide [39], and DALLE [44].
GenImage Dataset. This dataset mainly explores the Diffusion model for AI-generated image detection. Following its original setting [68], we adopt SD v1.4 as the training set and the remaining sets as the testing set, including Midjourney [37], SDv1.4 [45], SDv1.5 [45], ADM [14], GLIDE [39], Wukong [1], VQDM [19], and BigGAN [4]. As the varied image size of GenImage, following [55], we duplicate the images whose size is smaller than 224 and crop them to $224\times 224$ during training.
Metrics. Following the convention in C2P-CLIP[55], we use average precision (AP) and classification accuracy (Acc) for evaluation on the UniversalFakeDetect dataset, and only Acc on the GenImage dataset, with a 0.5 threshold for Acc across all datasets.
@@ 4.2 Implementation Details
Following [40], we adopt CLIP ViT-L/14 as the image encoder. In the training phase, we choose Adam [2] as the optimizer with the learning rate of $5\times 10^{-5}$. The batch size is set to 16 on UniversalFakeDetect and 32 on GenImage. The model is trained on a single NVIDIA 3090 GPU with only one epoch. The feature dimension $D$ of the image encoder is set to 1024. $N_{a}$ is set to 6 and $N_{t}$ is set to 9. For the MLP-based adapter, the middle dimension $\hat{D}$ is set to 64, and the scale factor $s$ is set to 0.1. For the Conditional Information Learner, the patch number $N_{p}$ is set to 192. The number of learnable tokens $M$ inserted in each block is set to 2. For the Test-Time Token Tuning, the number of views $N_{v}$ is set to 32, the number of selected confident views $m$ is set to 6, the tuning step number $T$ is set to 2, and the learning rate during Test-Time Token Tuning is set to $5\times 10^{-3}$.
Table 1: Evaluation of cross-method performance on the UniversalFakeDetect dataset [40] using Accuracy (%). The results of RINE are cited from their original report. The results of AIDE are reproduced by its official pre-trained model. Results of the other methods are cited from C2P-CLIP. The best results are highlighted in bold, and the second-best results are underlined.
Methods GAN Deep fakes Low level Perceptual loss Guided LDM Glide Dalle mAcc
Pro- GAN Cycle- GAN Big- GAN Style- GAN Gau- GAN Star- GAN SITD SAN CRN IMLE 200 steps 200 w/cfg 100 steps 100 27 50 27 100 10
Spec [64] 49.90 99.90 50.50 49.90 50.30 99.70 50.10 50.00 48.00 50.60 50.10 50.90 50.40 50.40 50.30 51.70 51.40 50.40 50.00 55.45
F3Net [42] 99.38 76.38 65.33 92.56 58.10 100.00 63.48 54.17 47.26 51.47 51.47 69.20 68.15 75.35 68.80 81.65 83.25 83.05 66.30 71.33
UniFD [40] 100.00 98.50 94.50 82.00 99.50 97.00 66.60 63.00 57.50 59.50 72.00 70.03 94.19 73.76 94.36 79.07 79.85 78.14 86.78 81.38
LGrad [52] 99.84 85.39 82.88 94.83 72.45 99.62 58.00 62.50 50.00 50.74 50.78 77.50 94.20 95.85 94.80 87.40 90.70 89.55 88.35 80.28
FreqNet [53] 97.90 95.84 90.45 97.55 90.24 93.41 97.40 88.92 59.04 71.92 67.35 86.70 84.55 99.58 65.56 85.69 97.40 88.15 59.06 85.09
NPR [54] 99.84 95.00 87.55 96.23 86.57 99.75 76.89 66.94 98.63 50.00 50.00 84.55 97.65 98.00 98.20 96.25 97.15 97.35 87.15 87.56
FatFormer [35] 99.89 99.32 99.50 97.15 99.41 99.75 93.23 81.11 68.04 69.45 69.45 76.00 98.60 94.90 98.65 94.35 94.65 94.20 98.75 90.86
RINE [28] 100.00 99.30 99.60 88.90 99.80 99.50 80.60 90.60 68.30 89.20 90.60 76.10 98.30 88.20 98.60 88.90 92.60 90.70 95.00 91.31
AIDE  [61] 99.99 98.49 83.95 99.65 73.24 99.90 54.09 68.06 71.14 60.87 61.00 88.50 98.20 97.45 98.40 98.15 98.40 97.85 97.50 86.57
C2P-CLIP  [55] 99.98 97.31 99.12 96.44 99.17 99.60 93.77 95.56 64.38 93.29 93.29 69.10 99.25 97.25 99.30 95.25 95.25 96.10 98.55 93.79
@@
Table 2: Evaluation of cross-method performance on the UniversalFakeDetect dataset [40] using Average Precision (%). The results of previous methods are cited from the same sources as Table 1.
Methods GAN Deep fakes Low level Perceptual loss Guided LDM Glide Dalle mAP
Pro- GAN Cycle- GAN Big- GAN Style- GAN Gau- GAN Star- GAN SITD SAN CRN IMLE 200 steps 200 w/cfg 100 steps 100 27 50 27 100 10
Spec [64] 55.39 100.00 75.08 55.11 66.08 100.00 45.18 47.46 57.12 53.61 50.98 57.72 77.72 77.25 76.47 68.58 64.58 61.92 67.77 66.21
F3Net [42] 99.96 84.32 69.90 99.72 56.71 100.00 78.82 52.89 46.70 63.39 64.37 70.53 73.76 81.66 74.62 89.81 91.04 90.86 71.84 76.89
UniFD [40] 100.00 98.13 94.46 86.66 99.25 99.53 91.67 78.54 67.54 83.12 91.06 79.24 95.81 79.77 95.93 93.93 95.12 94.59 88.45 90.14
LGrad [52] 100.00 93.98 90.69 99.86 79.36 99.98 67.91 59.42 51.42 63.52 69.61 87.06 99.03 99.16 99.18 93.23 95.10 94.93 97.23 86.35
FreqNet [53] 99.92 99.63 96.05 99.89 99.71 98.63 99.92 94.42 74.59 80.10 75.70 96.27 96.06 100.00 62.34 99.80 99.78 96.39 77.78 91.95
NPR [54] 100.00 99.53 94.53 99.94 88.82 100.00 84.41 97.95 99.99 50.16 50.16 98.26 99.92 99.91 99.92 99.87 99.89 99.92 99.26 92.76
FatFormer [35] 100.00 100.00 99.98 99.75 100.00 100.00 97.99 97.94 81.21 99.84 99.93 91.99 99.81 99.09 99.87 99.13 99.41 99.20 99.82 98.16
RINE [28] 100.00 100.00 99.90 99.40 100.00 100.00 97.90 97.20 94.90 97.30 99.70 96.40 99.80 98.30 99.90 98.80 99.30 98.90 99.30 98.79
AIDE  [61] 100.00 99.89 94.44 100.00 97.69 100.00 76.35 83.60 88.82 90.73 99.94 97.28 99.29 99.12 99.34 99.30 99.25 99.19 99.03 95.96
C2P-CLIP  [55] 100.00 100.00 99.96 99.50 100.00 100.00 98.59 98.92 84.56 99.86 99.95 94.13 99.99 99.83 99.98 99.72 99.79 99.83 99.91 98.66
@@
Table 3: Evaluation of cross-method performance on the GenImage [68] dataset using Accuracy (%). The results of DRCT, CO-SPY, AIDE, and SAFE are cited for their original report. The other results are cited from C2P-CLIP.
@@
DeiT-S [56] ICML 2021 55.6 99.9 99.8 49.8 58.1 98.9 56.9 53.5 71.6
UnivFD [40] CVPR 2023 91.5 96.4 96.1 58.1 73.4 94.5 67.8 57.7 79.5
NPR  [54] CVPR 2024 81.0 98.2 97.9 76.9 89.8 96.9 84.1 84.2 88.6
FreqNet  [53] AAAI 2024 89.6 98.8 98.6 66.8 86.5 97.3 75.8 81.4 86.8
FatFormer [35] CVPR 2024 92.7 100.0 99.9 75.9 88.0 99.9 98.8 55.8 88.9
DRCT [6] ICML 2024 91.5 95.0 94.4 79.4 89.2 94.7 90.0 81.7 89.5
CO-SPY  [11] CVPR 2025 83.5 96.8 96.8 67.3 93.0 95.9 78.8 65.2 84.7
AIDE  [61] ICLR 2025 79.4 99.7 99.8 78.5 91.8 98.7 80.3 66.9 86.9
SAFE  [31] KDD 2025 95.3 99.4 99.3 82.1 96.3 98.2 96.3 97.8 95.6
C2P-CLIP  [55] AAAI 2025 88.2 90.9 97.9 96.4 99.0 98.8 96.5 98.7 95.8

### v3

@@ 3.3 Image-Adaptive Prompt Learning
Specifically, at the inference stage, the Test-Time Token Tuning starts with generating multiple views $V\in R^{N_{v}\times 224\times 224\times 3}$ from the original image, where $N_{v}$ is the number of views. In particular, the generated views could be divided into a global view and multiple local views. For a global view, we resize the entire image to the input size of CLIP, i.e. $224\times 224$. For local views, we randomly crop $N_{v}$-1 $224\times 224$ sub-images followed by horizontal flipping augmentation. If the image is too small to crop $N_{v}$-1 different sub-images, we randomly crop $N_{v}$-1 smaller square patches and then resize these patches to $224\times 224$. To reduce the noise from randomly generated views, the Confidence Selection is conducted. We select $m$ views with high predicted confidence for subsequent tuning. The predicted confidence $S_{c}^{i}$ is calculated via Eq. 2.
@@
where $z_{i}$ is the logit of the i-th view. $\sigma(.)$ is the Sigmoid function that transform $z_{i}$ into $(0,1)$.
Given the selected views, we tune the test-time adaptive tokens by minimizing entropy, while other parameters are frozen. The objective function mainly includes averaged entropy loss [shu2022test] and pointwise entropy loss [sheng2025r]. For the averaged entropy loss, its purpose is to improve consistency among predictions across the generated views, and it is shown in Eq. 3 and Eq. 4.
@@
Here, $z_{k}$ is the logit calculated by the model of the $k$-th selected view. Some methods [sheng2025r] adopt pointwise entropy loss to push the decision boundary away from low-entropy regions. Its objective function is shown in Eq. 5.
@@
where $p_{k}$ denotes $\sigma(z_{k})$ of k-th selected view. Based on the results in Tab. LABEL:tab:8, we ultimately adopt the averaged entropy loss due to its superior performance across all experimental datasets. Additionally, the above tuning process is performed for a total of $T$ steps.
Conditional Information Learner. To capture subtle and varying cues in AI-generated images, we propose the Conditional Information Learner, which extracts conditional information tailored to each input. The conditional information is partitioned into forgery-specific and general components, which act on the input prompt to enable the model to focus on the distinct characteristics of various inputs. Moreover, the condition is obtained from the most texture-rich region with a high-pass filter, emphasizing low-level patterns often overlooked by the high-level semantic features of the pre-trained CLIP backbone.
Specifically, for a given input image, we divide it into local patches $P_{l}\in R^{N_{p}\times 32\times 32\times 3}$, where $N_{p}$ is the number of patches. We use the DCT score [yan2024sanity] to measure the texture richness of each patch and select the patch with the highest DCT value for conditional information extraction. After that, the selected patches are input into high-pass filters [fridrich2012rich] to obtain the high-frequency pattern. Two CNN-based feature extractors (see the supplementary for details) with identical architectures but separate parameters are used to extract the forgery-specific condition $C_{f}$ and the general condition $C_{g}$, respectively. An auxiliary supervision branch is attached after a forgery-specific condition. It consists of a single-layer MLP, enabling the supervised branch to extract features strongly correlated with the presence of forgery, while the unsupervised branch learns more general features related to the image status. Moreover, as the conditional information is extracted from a small patch, this process is both time-efficient and memory-efficient.
After acquiring the conditional information $C_{g}$ and $C_{f}$, we fuse them with the test-time adaptive tokens $A\in\mathbb{R}^{2\times D}$ via the learnable scaling factor, to obtain the image-adaptive prompt $P\in\mathbb{R}^{2\times D}$ as shown in Eq. 6.
$P=\{\alpha_{f}*C_{f}+A[0,:],\alpha_{g}*C_{g}+A[1,:]\},$ (6)
where $\alpha_{f}$ and $\alpha_{g}$ denote the learnable coefficients associated with the forgery-specific and general conditional information, respectively, and $\{\,\cdot\,$$\cdot\,\}$ represents the concatenation operation.
@@
3.4 Optimal Input Selection
Since images generated by different models often vary considerably in resolution and aspect ratio, previous methods typically adopt resizing or center cropping to standardize the input. However, such fixed preprocessing strategies may discard or distort critical artifacts, as the most discriminative regions for forgery detection differ across images. To address this limitation, we propose Optimal Input Selection, which only occurs during inference. Specifically, we first generate diverse inputs for a single test image. After testing-time token tuning, these inputs are fed into the tuned model to yield predictions. Their confidence scores are computed using the same procedure as Eq. 2, and the prediction with the highest score is taken as the final decision. Additionally, we reuse the $m$ views selected through confidence selection as the inputs for convenience.
@@ 4.1 Datasets and Metrics
We conduct experiments on two widely used datasets, i.e., UniversalFakeDetect Dataset [ojha2023towards] and GenImage Dataset [zhu2023genImage], to demonstrate the effectiveness of our framework.
UniversalFakeDetect Dataset. This dataset includes 20 subsets of generated images. Following previous protocol [ojha2023towards, wang2020cnn], the real and fake images in ProGAN [progan] are used for training, while the testing set contains 19 subsets from ProGAN and other 18 different generative models, including CycleGAN [Zhu-2017-cycleGAN], BigGAN [brock2018brock], StyleGAN [stylegan], GauGAN [park2019SPADE], StarGAN [choi2018stargan], DeepFakes [roessler2019faceforensicspp], SITD [sitd], SAN [san], CRN [crn], IMLE [imle], Guided [dhariwal2021diffusion], LDM [rombach2022high], Glide [glide], and DALLE [dalle-orig].
GenImage Dataset. This dataset mainly explores the Diffusion model for AI-generated image detection. Following its original setting [zhu2023genImage], we adopt SD v1.4 as the training set and the remaining sets as the testing set, including Midjourney [Midjourney], SDv1.4 [rombach2022high], SDv1.5 [rombach2022high], ADM [dhariwal2021diffusion], GLIDE [glide], Wukong [Wukong], VQDM [gu2022vector], and BigGAN [brock2018brock]. As the varied image size of GenImage, following [tan2025c2p], we duplicate the images whose size is smaller than 224 and crop them to $224\times 224$ during training.
Metrics. Following the convention in C2P-CLIP[tan2025c2p], we use average precision (AP) and classification accuracy (Acc) for evaluation on the UniversalFakeDetect dataset, and only Acc on the GenImage dataset, with a 0.5 threshold for Acc across all datasets.
@@ 4.2 Implementation Details
Following [ojha2023towards], we adopt CLIP ViT-L/14 as the image encoder. In the training phase, we choose Adam [adam2014method] as the optimizer with the learning rate of $5\times 10^{-5}$. The batch size is set to 16 on UniversalFakeDetect and 32 on GenImage. The model is trained on a single NVIDIA 3090 GPU with only one epoch. The feature dimension $D$ of the image encoder is set to 1024. $N_{a}$ is set to 6 and $N_{t}$ is set to 9. For the MLP-based adapter, the middle dimension $\hat{D}$ is set to 64, and the scale factor $s$ is set to 0.1. For the Conditional Information Learner, the patch number $N_{p}$ is set to 192. The number of learnable tokens $M$ inserted in each block is set to 2. For the Test-Time Token Tuning, the number of views $N_{v}$ is set to 32, the number of selected confident views $m$ is set to 6, the tuning step number $T$ is set to 2, and the learning rate during Test-Time Token Tuning is set to $5\times 10^{-3}$.
Table 1: Evaluation on the UniversalFakeDetect dataset [ojha2023towards] using Accuracy (%). The results of RINE are cited from their original report. The results of AIDE are reproduced by its official pre-trained model. Results of the other methods are cited from C2P-CLIP. $\clubsuit$: evaluating based on the official pre-trained model. The 4-class ProGAN is used as the training data.
Methods GAN
Deep
fakes
Low level Perceptual loss Guided LDM Glide Dalle mAcc
Pro-
GAN
Cycle-
GAN
Big-
GAN
Style-
GAN
Gau-
GAN
Star-
GAN
SITD
SAN
CRN
IMLE
200
steps
200
w/cfg
100
steps
100
27
50
27
100
10
UniFD [ojha2023towards] 100.00 98.50 94.50 82.00 99.50 97.00 66.60 63.00 57.50 59.50 72.00 70.03 94.19 73.76 94.36 79.07 79.85 78.14 86.78 81.38
LGrad [tan2023learning] 99.84 85.39 82.88 94.83 72.45 99.62 58.00 62.50 50.00 50.74 50.78 77.50 94.20 95.85 94.80 87.40 90.70 89.55 88.35 80.28
FreqNet [tan2024frequency] 97.90 95.84 90.45 97.55 90.24 93.41 97.40 88.92 59.04 71.92 67.35 86.70 84.55 99.58 65.56 85.69 97.40 88.15 59.06 85.09
NPR [tan2024rethinking] 99.84 95.00 87.55 96.23 86.57 99.75 76.89 66.94 98.63 50.00 50.00 84.55 97.65 98.00 98.20 96.25 97.15 97.35 87.15 87.56
FatFormer [liu2024forgery] 99.89 99.32 99.50 97.15 99.41 99.75 93.23 81.11 68.04 69.45 69.45 76.00 98.60 94.90 98.65 94.35 94.65 94.20 98.75 90.86
RINE [koutlis2024leveraging] 100.00 99.30 99.60 88.90 99.80 99.50 80.60 90.60 68.30 89.20 90.60 76.10 98.30 88.20 98.60 88.90 92.60 90.70 95.00 91.31
AIDE  [yan2024sanity] 99.99 98.49 83.95 99.65 73.24 99.90 54.09 68.06 71.14 60.87 61.00 88.50 98.20 97.45 98.40 98.15 98.40 97.85 97.50 86.57
C2P-CLIP  [tan2025c2p] 99.98 97.31 99.12 96.44 99.17 99.60 93.77 95.56 64.38 93.29 93.29 69.10 99.25 97.25 99.30 95.25 95.25 96.10 98.55 93.79
SAFE [li2025improving] 99.86 98.87 89.73 98.04 91.52 99.90 93.10 85.56 95.91 50.10 50.10 82.35 98.80 98.65 98.80 95.75 96.60 97.25 97.50 90.44
MiraGe [shi2025mirage] 100.00 94.30 96.50 96.80 93.60 96.10 88.70 75.80 71.90 92.90 92.90 82.00 98.30 94.60 98.60 97.50 97.50 98.00 98.60 92.90
@@
Table 2: Evaluation on the UniversalFakeDetect dataset [ojha2023towards] using Average Precision (%). The results of previous methods are cited from the same sources as Table 1. The 4-class ProGAN is used as the training data.
Methods GAN
Deep
fakes
Low level Perceptual loss Guided LDM Glide Dalle mAP
Pro-
GAN
Cycle-
GAN
Big-
GAN
Style-
GAN
Gau-
GAN
Star-
GAN
SITD
SAN
CRN
IMLE
200
steps
200
w/cfg
100
steps
100
27
50
27
100
10
UniFD [ojha2023towards] 100.00 98.13 94.46 86.66 99.25 99.53 91.67 78.54 67.54 83.12 91.06 79.24 95.81 79.77 95.93 93.93 95.12 94.59 88.45 90.14
LGrad [tan2023learning] 100.00 93.98 90.69 99.86 79.36 99.98 67.91 59.42 51.42 63.52 69.61 87.06 99.03 99.16 99.18 93.23 95.10 94.93 97.23 86.35
FreqNet [tan2024frequency] 99.92 99.63 96.05 99.89 99.71 98.63 99.92 94.42 74.59 80.10 75.70 96.27 96.06 100.00 62.34 99.80 99.78 96.39 77.78 91.95
NPR [tan2024rethinking] 100.00 99.53 94.53 99.94 88.82 100.00 84.41 97.95 99.99 50.16 50.16 98.26 99.92 99.91 99.92 99.87 99.89 99.92 99.26 92.76
FatFormer [liu2024forgery] 100.00 100.00 99.98 99.75 100.00 100.00 97.99 97.94 81.21 99.84 99.93 91.99 99.81 99.09 99.87 99.13 99.41 99.20 99.82 98.16
RINE [koutlis2024leveraging] 100.00 100.00 99.90 99.40 100.00 100.00 97.90 97.20 94.90 97.30 99.70 96.40 99.80 98.30 99.90 98.80 99.30 98.90 99.30 98.79
AIDE  [yan2024sanity] 100.00 99.89 94.44 100.00 97.69 100.00 76.35 83.60 88.82 90.73 99.94 97.28 99.29 99.12 99.34 99.30 99.25 99.19 99.03 95.96
C2P-CLIP  [tan2025c2p] 100.00 100.00 99.96 99.50 100.00 100.00 98.59 98.92 84.56 99.86 99.95 94.13 99.99 99.83 99.98 99.72 99.79 99.83 99.91 98.66
SAFE [li2025improving] 100.00 99.77 95.88 99.93 97.19 100.00 97.45 85.07 99.27 41.32 47.23 95.81 99.96 99.92 99.96 98.93 99.21 99.36 99.67 92.42
MiraGe [shi2025mirage] 100.00 100.00 99.90 99.80 99.90 99.90 96.00 93.90 84.70 99.90 100.00 96.40 99.90 99.10 99.90 99.80 99.70 99.80 99.90 98.30
@@
Table 3: Evaluation on the GenImage [zhu2023genImage] dataset using Accuracy (%). The results of AIDE, DRCT, and ATTSD are cited for their original reports. $\spadesuit$: re-implementing on SDv1.4 setting. The others are cited from C2P-CLIP. The SDv1.4 is used as the training set.
@@
UniFD [ojha2023towards] CVPR 2023 91.5 96.4 96.1 58.1 73.4 94.5 67.8 57.7 79.5
NPR  [tan2024rethinking] CVPR 2024 81.0 98.2 97.9 76.9 89.8 96.9 84.1 84.2 88.6
FreqNet  [tan2024frequency] AAAI 2024 89.6 98.8 98.6 66.8 86.5 97.3 75.8 81.4 86.8
FatFormer [liu2024forgery] CVPR 2024 92.7 100.0 99.9 75.9 88.0 99.9 98.8 55.8 88.9
DRCT [chendrct] ICML 2024 91.5 95.0 94.4 79.4 89.2 94.7 90.0 81.7 89.5
AIDE  [yan2024sanity] ICLR 2025 79.4 99.7 99.8 78.5 91.8 98.7 80.3 66.9 86.9
C2P-CLIP  [tan2025c2p] AAAI 2025 88.2 90.9 97.9 96.4 99.0 98.8 96.5 98.7 95.8
SAFE  [li2025improving] KDD 2025 95.7 99.9 99.8 59.5 91.7 95.7 98.0 77.4 90.3
ATTSD  [cai2025adaptive] ICCV 2025 91.8 94.7 93.8 85.0 97.3 91.4 81.2 90.2 90.7
MiraGe [shi2025mirage] ACMMM 2025 83.2 98.8 98.5 82.7 91.3 97.6 92.4 96.5 92.6